OpenStudy (anonymous):

Simplifying Expressions Using the Distributive Property ; -2(n-5)

OpenStudy (anonymous):

Just told you the logic.......in previous question -2(n - 5) = -2*n - (-2)*5 = -2n + 10

OpenStudy (anonymous):

-2n+10